# AllSpark[Uncaptioned image]: A Multimodal Spatio-Temporal General Intelligence Model with Thirteen Modalities

### v2

@@ IV-B Dataset
Table 4 summarizes the data and tasks used in the experiments for each modality. Below, we will provide detailed explanations in order:
Text: The IMDB[43] dataset is a binary sentiment analysis dataset consisting of 50,000 reviews from the Internet Movie Database (IMDb) labelled as positive or negative. Additionally, the dataset includes some unlabelled data. In our experiments, only the labelled data from the IMDB dataset were utilized for supervised sentiment classification tasks.
@@
Infrared: SYSU-MM01[51] is a cross-modal human reidentification dataset with optical-infrared modalities that includes images from 4 RGB cameras and 2 infrared cameras of 491 different individuals. We adopted the official dataset split, with the training set containing 20,284 RGB images and 9,929 infrared images from 296 individuals and the test set consisting of 3,803 infrared images from 96 individuals used as queries and 301 randomly selected RGB images for reidentification targets.
@@
Oblique Photography: WHU-OMVS[42] is an oblique photography dataset designed for 3D reconstruction tasks. The dataset provides imagery from five different viewpoints, along with camera parameters and other relevant information. It consists of six areas, and in our experiments, area 1 is used as the training set, and area 2 is used as the test set.
Video: The UCF101[52] dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips. The videos have a combined duration of 27 hours and a resolution of 320*240 pixels and were sourced from YouTube.
@@
IV-C Results
@@ IV-C1 RGB
TABLE III: rgb image classification with allspark
Method
Publication
Acc(%)
CNN-CapsNet[54]
RS2019
89.03
DFAGCN[55]
TNNLS2021
89.29
D-CNN with GoogleNet[56]
TGRS2018
90.49
D-CNN with VGGNet[56]
TGRS2018
91.89
SCCov[57]
TNNLS2019
92.1
SeCo-ResNet-50[58]
ICCV2021
92.91
MG-CAP[59]
TIP2020
92.95
LSENet[60]
TIP2021
93.34
MSANet[61]
JSTARS2021
93.52
IDCCP[62]
TGRS2021
93.76
MBLANet[63]
TIP2021
94.66
GRMANet-ResNet-50[64]
TGRS2021
94.72
EMSNet[65]
TGRS2023
95.37
ViTAE-B + RVSA[66]
TGRS2022
95.69
AllSpark
ours
94.85
@@
We evaluated the performance of AllSpark on the RGB image scene classification task using the NWPU-RESISC45 dataset, with the top-1 accuracy as the evaluation metric. AllSpark leverages expert knowledge from Lynx by loading its pretrained weights. Therefore, in Table III, we compare AllSpark with the state-of-the-art models known on the NWPU-RESISC45 dataset. The results indicate that AllSpark outperforms most baseline models, with a margin of only 0.84 compared to that of the SOTA (95.69). This highlights AllSpark’s exceptional perception and interpretation capabilities in the RGB modality.
IV-C2 MSI
@@
We evaluated the performance of AllSpark on the MSI scene classification task using the EuroSAT dataset. In the experiment, all 13 spectral bands of the images were simultaneously input into the model. The model’s objective was to correctly classify the images into one of the 10 specified categories, and the evaluation metric chosen was the top-1 accuracy. AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups, as shown in Table IV: those with expert knowledge intervention and those without. Expert knowledge intervention refers to baseline models pretraining on large datasets such as BigEarthNet and then fine-tuning on the EuroSAT dataset, while no expert knowledge indicates baseline models trained directly from scratch on the EuroSAT dataset. The results show that our model outperforms most models in the no expert knowledge group, with a margin of only 2.60 compared to the state-of-the-art model (ResNet-152). Furthermore, AllSpark lags behind the best result in the expert knowledge group by only 4.75, demonstrating its excellent adaptability to the multispectral modality.
TABLE IV: msi land cover classification with allspark
Expert Know.
Method
Publication
Acc(%)
No
ResNet-18[58]
ICCV2021
63.21
ResNet-50[67]
IGARSS2023
91.13
InceptionNet[68]
ICIP2020
93.07
EfficientNet[69]
IJSTARS2021
93.94
@@
ResNet-152[67]
IGARSS2023
96.63
Yes
MoCoV2[58]
ICCV2021
89.51
SeCo[58]
ICCV2021
93.14
SEER[70]
Arxiv2022
97.6
DINO-MC[71]
Arxiv2023
98.78
IV-C3 HSI
We conducted a pixel classification task on the Pavia University dataset for the hyperspectral modality. The model treats all spectral bands of a single pixel as one sample and predicts the land cover category of that pixel. The reported metrics include overall accuracy (OA), average accuracy (AA), and kappa. Since AllSpark does not possess expert knowledge of the hyperspectral modality, we compared it with the semisupervised baselines summarized by D. Uchaev and D. Uchaev [72]. The results in Table V demonstrate that AllSpark outperforms many hyperspectral image classification methods, such as IFRF and S-DMM, by a factor of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
TABLE V: hsi pixel classification with allspark
Method
OA
AA
Kappa
@@ IV-C4 Table
TABLE VI: PM2.5 prediction with allspark
@@
Yes Stacked ResNet-LSTM[74] 40.68 23.75 0.8
@@
Table VI presents the comparative results between AllSpark and the baselines. It is worth noting that some works specifically focus on the prediction task on the PRSA dataset and design expert models with specific architectures based on dataset characteristics. For example, the CBAM-CNN-Bi-LSTM proposed by D. Li et al.[75] used a CNN to extract spatial dependencies between air monitoring stations and Bi-LSTM to capture the temporal dependencies of PM2.5 data. Similarly, the stacked ResNet-LSTM model proposed by X. Cheng et al.[74] employs a stacking LSTM strategy to enhance the extraction of temporal features in PM2.5 data. Therefore, we categorize baseline methods into two types: those with an expert architecture and those without an expert architecture. Among the models without expert architecture, our approach achieves the best performance among the baselines and is slightly inferior to the state-of-the-art method for models with expert architecture (CBAM-CNN-Bi-LSTM). This reflects the excellent adaptability of AllSpark to the table modality.
IV-C5 Code
@@
TABLE VII: code document generation with allspark
@@
Like in the previous modalities, since MSI-AGI does not possess expert knowledge of the code modality, we categorized the baselines into two groups: those with and without expert knowledge. As shown in Table VII, AllSpark achieved SOTA results in the group without expert knowledge, and the results were comparable to those of models trained with expert knowledge. This finding demonstrates the strong adaptability of AllSpark to the code modality.
IV-C6 PointCloud
@@
TABLE VIII: pointcloud classification with allspark
Method
Expert Architecture Expert Knowledge Acc(%)
PointNet[1]
Yes No 89.2
Kd-net[77]
Yes No 90.6
SPH3D-GCN[78]
Yes No 91.4
PointNet++[79]
Yes No 91.9
SO-Net[80]
Yes No 92.5
PointVGG[81]
Yes No 93.6
PointBERT[41]
No Yes 93.8
PointGPT[82]
Yes Yes 94.9
AllSpark
No No 91.2
The AllSpark module is designed based on a general sequence-to-sequence architecture. As shown in Table VIII, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-Net). This approach maintains comparability with the state-of-the-art PointGPT model, which has both a modal expert structure and modal expert knowledge. This finding suggested that AllSpark has significant potential for applications in the point cloud modality.
IV-C7 Trajectory
@@
In Table IX, AllSpark is compared with the state-of-the-art trajectory prediction models. The results indicate that AllSpark, which uses a unified structure without trajectory modality expert knowledge, outperforms most expert models. It achieves a prediction accuracy close to that of the SOTA model (STAR), with a difference of only 0.07 in the ADE metric and 0.11 in the FDE metric. This finding suggested that AllSpark demonstrated excellent adaptability to the trajectory modality.
TABLE IX: trajectory prediction with allspark
Method
ADE/FDE
Social GAN
0.87/1.62
SoPhie
0.70/1.43
STAR
0.36/0.64
SGCN
0.63/1.03
CAGN
0.41/0.65
SIT
0.39/0.62
SocialVAE
0.47/0.76
PCENet
0.54/0.87
AgentFormer
0.45/0.75
MemoNet
0.40/0.61
SocialVAE+FPC
0.41/0.58
TUTR
0.40/0.61
AllSpark
0.43/0.69
IV-C8 SAR
@@
For the SAR modality, the adaptability of AllSpark was tested on the MSTAR dataset, where the model is required to identify SAR images of ten military targets, and the metric used is the top-1 accuracy. In the experiment, the preprocessing of the MSTAR dataset followed the SOC settings from AConvNets[22]. Table X presents the comparison results between AllSpark and the state-of-the-art model under these settings. Since AllSpark lacks expert knowledge of SAR image modalities and spatial invariance prior to changes in perspective; its performance on the MSTAR dataset is suboptimal, with an accuracy of 88%. This approach is comparable to the EMACH method. Enhancing the model’s adaptability to SAR images and improving its ability to extract spatial information are directions for future research.
TABLE X: sar classification with allspark
Method
Acc(%)
EMACH[83]
88
SVM[83]
90
AdaBoost[83]
92
MSRC[84]
93.6
IGT[83]
95
MSS[85]
96.6
Cond Gauss[86]
97
M-PMC[87]
98.8
AConvNets[22]
99.13
AllSpark
88
IV-C9 Infrared
AllSpark’s adaptability to the infrared modality was tested on the SYSU-MM01 dataset in a cross-modal human reidentification task involving RGB and infrared modalities. In this experiment, the model receives both an infrared image and an RGB image and is tasked with identifying and matching the same person across different modality images. Table XI compares AllSpark and the other state-of-the-art reidentification methods, with the evaluation metric being rank at an accuracy of 20. The results indicate that AllSpark (76.31%) achieved similar accuracy to that of HSME (77.95%). While our method did not reach the highest accuracy, it surpassed some expert models in RGB-infrared reidentification (e.g., TONE, HCML) when using a unified architecture without modal expert knowledge, demonstrating the potential of AllSpark in handling the infrared modality.
TABLE XI: visible-infrared reidentification with allspark
Method
Rank-20(All-search)
Two-stream[51]
65.5
One-stream[51]
66.74
TONE[88]
68.6
HCML[88]
69.17
Zero-Pad[51]
71.33
AllSpark 76.31
HSME[89]
77.95
BDTR[90]
81.07
DDAG[38]
95.81
IV-C10 Graph
TABLE XII: traffic prediction with allspark
Method
RMSE MAE R2
HI
6.8 14.2 10.15
GWNet
3.51 7.28 9.96
DCRNN
3.54 7.47 10.32
AGCRN
3.59 7.45 10.47
STGCN
3.6 7.43 10.35
GTS
3.59 7.44 10.25
MTGNN
3.47 7.21 9.7
STNorm
3.57 7.51 10.24
GMAN
3.44 7.35 10.07
PDFormer
3.62 7.47 10.91
STID
3.55 7.55 10.95
STAEformer
3.34 7.02 9.7
AllSpark
3.81 7.52 11.24
For the graph modality, the performance of AllSpark was evaluated on the traffic flow prediction task using the METR-LA dataset. The evaluation metrics include the RMSE, MAE, and R2. Table XII compares AllSpark and the state-of-the-art methods on the METR-LA dataset, with the baseline derived from[39]. The results show that AllSpark, without the intervention of modal expert knowledge, is only 0.47 away from the best result in terms of the RMSE, demonstrating its excellent adaptability to the graph modality.
IV-C11 Language
We tested AllSpark’s natural language processing capabilities on the IMDB dataset[43] with the task of binary sentiment classification (positive or negative). AllSpark loads weights from the Lynx; therefore, it can be considered to possess expert knowledge in the natural language modality. Compared with the SOTA models on the IMDB dataset, as shown in Table XIII, AllSpark outperformed most language models, with only a 0.32 difference from the SOTA result. This highlights AllSpark’s powerful understanding and analysis capabilities in natural language.
TABLE XIII: text understanding with allspark
Method
Acc(%)
RoBERTa[91]
95.3
ULMFiT[92]
95.4
BERT[93]
95.49
Mixed VAT[94, 95]
95.68
LongFormer[96]
95.7
XLNet[97]
96.8
ERNIE-Doc-Large[98]
97.1
AllSpark
96.78
IV-C12 Video
For the video modality, we evaluated AllSpark’s performance on action recognition tasks using the UCF101 dataset. The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being the top-1 accuracy. In Table XIV, we compare AllSpark with the current state-of-the-art models. Currently, AllSpark’s adaptation to the video modality is not optimal, as it shows a significant difference from the baseline model results. We attribute this to two main reasons: 1. The high redundancy in video information increases the training cost for AllSpark. We trained it for only 3 epochs on the UCF101 dataset. 2. AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective at capturing temporal information.
TABLE XIV: video classification with allspark
Method
Acc(%)
OPN
59.6
VCOP
72.4
SpeedNet
81.1
VTHCL
82.1
CVRL
94.4
VideoMAE v1
96.1
VideoMAE v2
99.6
AllSpark
27.5
IV-C13 Oblique Photography
For the oblique photography modality, we tested AllSpark’s performance on the 3D reconstruction task using the WHU-OMVS dataset. The model takes five-view images as input, and the goal is to output depth maps for reconstructing 3D models. The evaluation metric used was the percentage of accurate grids in total (PAG)[42], calculated by the following formula:
$PAG_{a}=\left(\frac{m_{a}}{m}\right)*100$ (7)
The suffixes in the PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 metres and ${PAG}_{10}$ indicates an error within 1 metre. Table XV compares AllSpark with popular multiview 3D reconstruction models on the WHU-OMVS dataset. AllSpark falls short in terms of accuracy compared to modality-specific expert models. We speculate two possible reasons: 1. Dense spatial information gradually diminishes in the deep structure of large models, a point verified in our exploratory experiments involving segmentation, detection, etc.; 2. The model architecture lacks flexibility and cannot connect gradual features like expert 3D reconstruction models such as Ada-MVS. Additionally, these methods lack the ability to design specific model structures for processing, severely restricting their performance. Adapting multimodal large models to dense prediction tasks and optimizing the architecture are future research directions.
TABLE XV: 3D reconstruction with allspark
Method
${PAG}_{6}$
${PAG}_{10}$
MVSNet
81.15 91.44
CasMVSNet
95.45 98.02
Ada-MVS
96.14 98.1
UCSNet
96.25 98.45
AllSpark
6.4 10.4

### v3

@@ IV-B Dataset
Below, we will provide detailed explanations of datasets in order:
Language: The IMDB[57] dataset is a binary sentiment analysis dataset consisting of 50,000 reviews from the Internet Movie Database (IMDb) labelled as positive or negative. Additionally, the dataset includes some unlabelled data. In our experiments, only the labelled data from the IMDB dataset were utilized for supervised sentiment classification tasks.
@@
TABLE IV: The few-shot classification results for the RGB modality
Model Training-free UC Merced RS19
5way-1shot 5way-5shot 5way-1shot 5way-5shot
MatchingNet[59] 48.18 67.39 67.68 85.01
RelationNet[60] 50.07 65.22 65.01 79.75
ProtoNet[61] 53.85 71.23 76.36 85.00
DLA-MatchNet[62] 53.76 63.01 68.27 79.89
SPNet[63] 57.64 73.52 81.06 88.04
AllSpark 95.58 97.64 97.16 98.94
TABLE V: The few-shot classification results for the point modality
Model Training-free ShapeNet ScanObject
5way-1shot 5way-5shot 5way-1shot 5way-5shot
FSLGNN[64] 64.98 76.14 29.91 32.77
RelationNet[60] 65.88 76.25 45.32 55.43
ProtoNet[61] 65.96 78.77 44.75 59.81
AllSpark 67.20 82.12 35.70 53.33
@@
IV-C Few-shot learning
To demonstrate AllSpark’s unique advantages over traditional models, we tested its few-shot performance on the RGB and point cloud modalities. It is worth noting that traditional few-shot learning methods typically require additional training steps. For instance, common approaches like ProtoNet[61] and MatchingNet[59] involve randomly splitting support and query sets on the training set for supervised training, a step known as meta-learning. Thanks to the integration of natural language, AllSpark requires no extra training step and can directly evaluate few-shot classification accuracy on the test set, significantly outperforming baseline models.
@@ IV-C1 RGB
We evaluated AllSpark’s RGB image few-shot classification performance on the UC-Merced[72] and WHU-RS19[73, 74] datasets, following the dataset splits from [62]. Specifically, the UC-Merced dataset uses six classes—Beach, Golf course, Mobile home park, River, Sparse residential, and Tennis court—as the test set, while the WHU-RS19 dataset uses five classes—Commercial, Meadow, Pond, River, and Viaduct—as the test set. AllSpark does not require meta-learning on the training set and is evaluated on 600 episodes directly on the test set, with 15 query samples per episode. As shown in the Table IV, we report the results for both 5-way 1-shot and 5-way 5-shot settings. The experiments demonstrate that AllSpark significantly outperforms baseline models without requiring any training, highlighting the advantage of LaRF.
IV-C2 Point
We evaluated AllSpark’s point cloud few-shot classification performance on the ShapeNet[75] and ScanObjectNN[76] datasets, following the dataset settings from [77]. Similar to the RGB modality, AllSpark does not require meta-learning on the training set and is evaluated on 700 episodes directly on the test set, with 15 query samples per episode. As shown in the Table V, we report the results for both 5-way 1-shot and 5-way 5-shot settings. The experiments demonstrate that AllSpark surpasses most baseline models even on previously unseen point cloud datasets, showcasing its outstanding semantic richness and generalization capabilities.
IV-D Ability to understand 10 spatio-temporal modalities
IV-D1 RGB
@@
We evaluated the performance of AllSpark on the RGB image scene classification task using the NWPU-RESISC45 dataset, with the top-1 accuracy as the evaluation metric. AllSpark leverages expert knowledge from Lynx by loading its pretrained weights. Therefore, in Table VI, we compare AllSpark with the state-of-the-art models. The results indicate that AllSpark outperforms most baseline models, with a margin of only 0.84 compared to that of the SOTA (95.69). This highlights AllSpark’s exceptional perception and interpretation capabilities in the RGB modality.
TABLE VI: RGB image classification with allspark
Method Publication Acc(%)
CNN-CapsNet[78] RS2019 89.03
DFAGCN[79] TNNLS2021 89.29
D-CNN with GoogleNet[80] TGRS2018 90.49
D-CNN with VGGNet[80] TGRS2018 91.89
SCCov[81] TNNLS2019 92.1
SeCo-ResNet-50[82] ICCV2021 92.91
MG-CAP[83] TIP2020 92.95
LSENet[84] TIP2021 93.34
MSANet[85] JSTARS2021 93.52
IDCCP[86] TGRS2021 93.76
MBLANet[87] TIP2021 94.66
GRMANet-ResNet-50[88] TGRS2021 94.72
EMSNet[89] TGRS2023 95.37
ViTAE-B + RVSA[90] TGRS2022 95.69
AllSpark ours 94.85
IV-D2 MSI
@@
We evaluated the performance of AllSpark on the MSI scene classification task using the EuroSAT dataset. In the experiment, all 13 spectral bands of the images were simultaneously input into the model. The model’s objective was to correctly classify the images into one of the 10 specified categories, and the evaluation metric chosen was the top-1 accuracy. AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups, as shown in Table VII: those with expert knowledge intervention and those without. Expert knowledge intervention refers to baseline models pretraining on large datasets such as BigEarthNet and then fine-tuning on the EuroSAT dataset, while no expert knowledge indicates baseline models trained directly from scratch on the EuroSAT dataset. The results show that our model outperforms most models in the no expert knowledge group, with a margin of only 2.60 compared to the state-of-the-art model (ResNet-152). Furthermore, AllSpark lags behind the best result in the expert knowledge group by only 4.75, demonstrating its excellent adaptability to the multispectral modality.
TABLE VII: MSI land cover classification with allspark
Expert Know. Method Publication Acc(%)
ResNet-18[82] ICCV2021 63.21
ResNet-50[91] IGARSS2023 91.13
InceptionNet[92] ICIP2020 93.07
EfficientNet[93] IJSTARS2021 93.94
@@
ResNet-152[91] IGARSS2023 96.63
MoCoV2[82] ICCV2021 89.51
SeCo[82] ICCV2021 93.14
SEER[94] Arxiv2022 97.6
DINO-MC[95] Arxiv2023 98.78
IV-D3 HSI
We conducted a pixel classification task on the Pavia University dataset for the hyperspectral modality. The model treats all spectral bands of a single pixel as one sample and predicts the land cover category of that pixel. The reported metrics include overall accuracy (OA), average accuracy (AA), and kappa. Since AllSpark does not possess expert knowledge of the hyperspectral modality, we compared it with the semisupervised baselines summarized by D. Uchaev and D. Uchaev [96]. The results in Table VIII demonstrate that AllSpark outperforms many hyperspectral image classification methods, such as IFRF and S-DMM, by a factor of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
TABLE VIII: HSI pixel classification with allspark
Method OA AA Kappa
@@ IV-D4 Table
TABLE IX: PM2.5 prediction with allspark
@@
Stacked ResNet-LSTM[98] 40.68 23.75 0.8
@@
Table IX presents the comparative results between AllSpark and the baselines. It is worth noting that some works specifically focus on the prediction task on the PRSA dataset and design expert models with specific architectures based on dataset characteristics. For example, the CBAM-CNN-Bi-LSTM proposed by D. Li et al.[99] used a CNN to extract spatial dependencies between air monitoring stations and Bi-LSTM to capture the temporal dependencies of PM2.5 data. Similarly, the stacked ResNet-LSTM model proposed by X. Cheng et al.[98] employs a stacking LSTM strategy to enhance the extraction of temporal features in PM2.5 data. Therefore, we categorize baseline methods into two types: those with an expert architecture and those without an expert architecture. Among the models without expert architecture, our approach achieves the best performance among the baselines and is slightly inferior to the state-of-the-art method for models with expert architecture (CBAM-CNN-Bi-LSTM). This reflects the excellent adaptability of AllSpark to the table modality.
IV-D5 Code
@@
TABLE X: code document generation with allspark
@@
Like in the previous modalities, since MSI-AGI does not possess expert knowledge of the code modality, we categorized the baselines into two groups: those with and without expert knowledge. As shown in Table X, AllSpark achieved SOTA results in the group without expert knowledge, and the results were comparable to those of models trained with expert knowledge. This finding demonstrates the strong adaptability of AllSpark to the code modality.
IV-D6 Point Cloud
@@
TABLE XI: pointcloud classification with allspark
Method Expert Architecture Expert Knowledge Acc(%)
PointNet[10] 89.2
Kd-net[101] 90.6
SPH3D-GCN[102] 91.4
PointNet++[103] 91.9
SO-Net[104] 92.5
PointVGG[105] 93.6
PointBERT[48] 93.8
PointGPT[106] 94.9
AllSpark 91.2
The AllSpark module is designed based on a general sequence-to-sequence architecture. As shown in Table XI, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-Net). This approach maintains comparability with the state-of-the-art PointGPT model, which has both a modal expert structure and modal expert knowledge. This finding suggested that AllSpark has significant potential for applications in the point cloud modality.
IV-D7 Trajectory
@@
In Table XII, AllSpark is compared with the state-of-the-art trajectory prediction models. The results indicate that AllSpark, which uses a unified structure without trajectory modality expert knowledge, outperforms most expert models. It achieves a prediction accuracy close to that of the SOTA model (STAR), with a difference of only 0.07 in the ADE metric and 0.11 in the FDE metric. This finding suggested that AllSpark demonstrated excellent adaptability to the trajectory modality.
TABLE XII: trajectory prediction with allspark
Method ADE/FDE
Social GAN 0.87/1.62
SoPhie 0.70/1.43
STAR 0.36/0.64
SGCN 0.63/1.03
CAGN 0.41/0.65
SIT 0.39/0.62
SocialVAE 0.47/0.76
PCENet 0.54/0.87
AgentFormer 0.45/0.75
MemoNet 0.40/0.61
SocialVAE+FPC 0.41/0.58
TUTR 0.40/0.61
AllSpark 0.43/0.69
IV-D8 SAR
@@
For the SAR modality, the adaptability of AllSpark was tested on the MSTAR dataset, where the model is required to identify SAR images of ten military targets, and the metric used is the top-1 accuracy. In the experiment, the preprocessing of the MSTAR dataset followed the SOC settings from AConvNets[34]. Table XIII presents the comparison results between AllSpark and the state-of-the-art model under these settings. AllSpark achieves 97.24% top1-accuracy, only 1.89% lower than the SOTA model. The experiments demonstrate that AllSpark is capable of effectively understanding SAR imagery.
TABLE XIII: sar classification with allspark
Method Acc(%)
EMACH[107] 88
SVM[107] 90
AdaBoost[107] 92
MSRC[108] 93.6
IGT[107] 95
MSS[109] 96.6
Cond Gauss[110] 97
M-PMC[111] 98.8
AConvNets[34] 99.13
AllSpark 97.24
IV-D9 Graph
For the graph modality, the performance of AllSpark was evaluated on the traffic flow prediction task using the METR-LA dataset. The evaluation metrics include the RMSE, MAE, and R2. Table XIV compares AllSpark and the state-of-the-art methods on the METR-LA dataset, with the baseline derived from[53]. The results show that AllSpark, without the intervention of modal expert knowledge, is only 0.47 away from the best result in terms of the RMSE, demonstrating its excellent adaptability to the graph modality.
IV-D10 Language
We tested AllSpark’s natural language processing capabilities on the IMDB dataset[57] with the task of binary sentiment classification (positive or negative). AllSpark loads weights from the Lynx; therefore, it can be considered to possess expert knowledge in the natural language modality. Compared with the SOTA models on the IMDB dataset, as shown in Table XV, AllSpark outperformed most language models, trailing the SOTA result by only 0.32. This highlights AllSpark’s powerful understanding and analysis capabilities in natural language.
TABLE XIV: traffic prediction with allspark
Method RMSE MAE R2
HI 6.8 14.2 10.15
GWNet 3.51 7.28 9.96
DCRNN 3.54 7.47 10.32
AGCRN 3.59 7.45 10.47
STGCN 3.6 7.43 10.35
GTS 3.59 7.44 10.25
MTGNN 3.47 7.21 9.7
STNorm 3.57 7.51 10.24
GMAN 3.44 7.35 10.07
PDFormer 3.62 7.47 10.91
STID 3.55 7.55 10.95
STAEformer 3.34 7.02 9.7
AllSpark 3.81 7.52 11.24
TABLE XV: text understanding with allspark
Method Acc(%)
RoBERTa[112] 95.3
ULMFiT[113] 95.4
BERT[114] 95.49
Mixed VAT[115, 116] 95.68
LongFormer[117] 95.7
XLNet[118] 96.8
ERNIE-Doc-Large[119] 97.1
AllSpark 96.78
IV-E Training and Inference Costs
As the scale of model parameters grows, the training and inference costs of large language models have been rapidly increasing. In this section, we provide detailed the training and inference costs of AllSpark for reference. All experiments are conducted on two NVIDIA A6000 48G GPUs, using a numerical precision of torch.float32. The hyperparameter settings are detailed in Section IV-A.
Due to the presence of modality-specific encoders, the activated parameters in AllSpark varies when processing different modalities. Table XVI summarizes the total parameters, trainable parameters, training time, and inference computational cost (measured in MACs) for each modality.
TABLE XVI: training and inference costs
Modality Total params.(M) Trainable params.(M) Training time Inference MACs(G)
RGB 8176.02 689.80 $\sim$18 hours 1695.98
MSI 7455.14 978.87 $\sim$30 hours 363.02
HSI 7450.58 974.31 $\sim$21 hours 374.61
Table 7486.67 1010.40 $\sim$44 hours 410.00
Code 6876.02 268.66 $\sim$59 hours 3606.88
Point 7594.81 1118.54 $\sim$49 hours 557.38
Trajectory 7381.00 904.73 $\sim$39 hours 538.58
SAR 7452.58 976.31 $\sim$87 hours 373.63
Graph 7223.91 747.64 $\sim$53 hours 1846.28
Language 6876.02 268.66 $\sim$34 hours 3606.88